OpenStudy (anonymous):

2a over 3 divided by one half. solve and show work. can you explain to me how to do this

OpenStudy (lollylau):

(2a/3)/(1/2) =2a/3 * 2 =4a/3

jhonyy9 (jhonyy9):

yes --- right

OpenStudy (anonymous):

Correct 4a/3.

OpenStudy (anonymous):

dividing by 1/2 is the same as multiplying by 2. Infact dividing by 1/x where \[x \neq 0\] is equivalent to multiplying by x because: \[\frac{1}{x} = x^{-1}\] so dividing by 1/x means multiplying by \[\frac{1}{\frac{1}{x}} = (x^{-1})^{-1} = x\] Thus \[\frac{2a}{3}\] divided by 1/2 is the same as \[\frac{4a}{3}\]

OpenStudy (lollylau):

do we have to use powers? (a/b)/(c/d)*(d/c)/(d/c) =(a/b)/(d/c).